# Sample, Don’t Search: Rethinking Test-Time Alignment for Language Models

### v1

@@ 3.3 Best-of-$n$ Sampling
This procedure implicitly biases the output distribution toward higher-reward responses. It is easy to see that in the limit of $n$, we are sampling from the maximum reward distribution; however, BoN is remarkably performant even when $n$ is small (Gao et al., 2022; Llama, 2024).
With some strong assumptions, we can show that the probability density of the BoN distribution can be approximated by the following form:
$p(y^{\ast}(n)\mid x)\propto\exp\!\left(\frac{r(x,y^{\ast}(n))}{\beta^{\ast}(n)}\right)\,p_{\text{LM}}(y^{\ast}(n)\mid x),$ (11)
where $\beta^{\ast}$ takes the form
$\beta^{\ast}(n)=\frac{\sigma_{d}(x)}{\sqrt{2\log n_{d}}-\frac{(\log\log n_{d}+\log(4\pi))}{2\sqrt{2\log n_{d}}}}.$ (12)
where $\sigma_{d}(x)$ is the standard deviation of the dominant component of the distribution of reward, under the assumption of Gaussian mixture, which we empirically observe (see Appendix C), when evaluated on samples from our base LM conditioned on $x$, and $n_{d}=w_{d}(x)n$, the dominant component’s sample size, where $w_{d}(x)$ is the weight of the dominant component. A full derivation of this approximation is provided in Appendix A. Note that this density equals our target distribution $\pi(y\mid x)$ from Eq. 2 if we set $\beta=\beta^{\ast}(n)$.
This rough equivalence, not noted in past literature to our knowledge, provides a new insight into a key distinction between BoN and the other methods we have discussed, including QAlign. In QAlign, we explicitly define a target distribution $\pi(y\mid x)$, and increasing the test-time budget $n$ refines our approximation of this fixed objective. In contrast, BoN does not converge to a predefined distribution; instead, increasing $n$ progressively shifts the selection process toward responses with higher rewards and deviating further from $p_{\text{LM}}(y\mid x)$. Insofar as the reward model is well-aligned with human preferences, this should make BoN highly effective. However, in realistic settings where the reward is imperfect, increasing $n$ can lead to over-optimization, i.e., selecting responses that score well under the reward model but degrade in actual quality. As a result, BoN can suffer from diminishing returns or even performance degradation as test-time compute increases, which we observe in our experiments (§4.14.2) and previous work (Gao et al., 2022).
@@ 4.1 Task-Specific Tuning
We evaluate QAlign for task-specific tuning using Llama-3.1-8B-Instruct as the base model and a custom Bradley & Terry (1952) RM trained on our preference dataset for mathematical reasoning. This RM was initialized with Llama-3.1-8B-Instruct and trained on 64 model-generated response pairs per prompt. We selected pairs based on ground truth answers. [^2] Data and RM will be provided upon publication.
@@
We compare against BoN, and WMV sampling applied to Llama-3.1-8B-Instruct using our trained RM and MV from samples from Llama-3.1-8B-Instruct. For each of the methods we generate $4096$ solutions per problem.
@@
We evaluate on GSM8K (Cobbe et al., 2021), a dataset of grade school math problems, and GSM-Symbolic (Mirzadeh et al., 2024), a new benchmark designed to assess out-of-distribution generalization in mathematical reasoning. As demonstrated by Mirzadeh et al. (2024), GSM-Symbolic exposes substantial performance degradation across state-of-the-art LLMs exhibiting drops of up to $65\%$ compared to their GSM8K scores when faced with simple variations like changed numerical values and extra redundant clauses.
@@ Results
The results of the task-specific tuning experiments are summarized in Figure 2. QAlign shows progressively lower error rates across both problems as we spend more computational resources. In contrast, on GSM8K, MV shows initial improvement but quickly saturates, while BoN displays improvement with additional compute, temporarily outperforming QAlign, but eventually reaches an inflection point, after which error rates begin to increase. This behavior aligns with the observations from Gao et al. (2022) and theoretical results from §3.3. WMV performs well on GSM8K but becomes relatively worse after a budget of approximately $2^{8}\times 10^{12.56}$ FLOPs.
While all methods show some performance drop on GSM-Symbolic compared to GSM8K, the magnitude varies significantly. BoN and WMV, which rely on the RM, struggle more with the distribution shift, likely because the RM is fit to particular GSM8K-specific patterns. Relative to these two, MV shows greater robustness to these changes. This is primarily because under the distribution shift the RM becomes less reliable, and since MV does not use it, MV’s performance degrades less. However, even under these circumstances, QAlign is able to leverage the RM and match MV on this dataset with enough compute. This suggests that QAlign is robust to imperfections in the RM and able to extract useful signal even when the RM’s reliability is compromised.
@@ 4.2 General Alignment
We evaluate QAlign for more general alignment using Tülu3-8B-SFT as the base model and Tülu3-8B-RM as the RM. The Tülu3 model family (Lambert et al., 2024) was selected for being fully open source, providing instruction-tuned, and aligned versions, along with their corresponding training code and datasets. This enables direct comparison of QAlign on top of the instruction-tuned model to the aligned model, sharing the initial starting point. Additionally, they provide an RM trained on the data used to train the DPO model, which we use in our experiments.
@@
We evaluate again on GSM8K; we add MATH500, a verified subset of the MATH test dataset that contains advanced mathematical problems across algebra, geometry, and calculus (Hendrycks et al., 2021), MMLU-Redux (Gema et al., 2025), a refined subset of commonsense and academic knowledge questions, TruthfulQA (Lin et al., 2022), which contains misleading questions designed to elicit truthful responses, and IFEval (Zhou et al., 2023b), which measures adherence to complex multi-step instructions. Among these datasets, IFEval is an open-ended generation and, therefore, requires the use of MBR with ROUGE-$1$ instead of MV and WMV. For simplicity, we still refer to the method as MV and WMV to denote MBR from base model samples and the weighted version with importance sampling, respectively, in our comparisons.
@@
For all datasets, we sampled the model with a temperature of 1.0 and prompted it to provide reasoning. Complete prompts are available in Appendix D. As with the mathematical reasoning experiments, we tuned the $\beta$ parameter (in this case, $\beta=0.5$), for Tülu3-8B-RM, using $128$ samples from the GSM8K training data to achieve an average acceptance rate of $50\%$, and using the same $\beta$ for all datasets.
@@ Results
The results of the general alignment experiments are summarized in Table 1. Figure 1 plots the average error rate across all of the datasets as a function of the floating point operations (FLOPS), and Appendix E contains all of the error plots for each individual problem. Similar to our task-specific findings, QAlign, MV, and WMV consistently reduce error rates across all five general alignment datasets as computation increases, with MV exhibiting early saturation. However, similar to the results from GSM-Symbolic, MV outperforms both WMV and BoN. MV applied to the DPO model begins with lower error rates but saturates earlier, ultimately reaching error rates that are close to but still worse than QAlign’s final results. Furthermore, we observe that BoN’s inflection point occurs at a lower computational budget than observed in task-specific experiments. This suggests that for real-world alignment problems, BoN is not an effective test-time approach, while QAlign maintains its improvement trajectory even with general RMs.
@@
Table 1: Overview of the results on general alignment. Results show accuracy percentages with fixed 1024 sampled solutions across different methods and models. Our proposed QAlign outperforms other approaches on most benchmarks, achieving the highest scores on 5 out of 5 datasets when applied to Tülu3-8B-SFT, and attaining an average performance better than Tülu3-8B-DPO. The highlighted cells indicate best result per benchmark.
@@ 5 Related Work
Our work builds on Quest (Faria et al., 2024), which uses MCMC to sample diverse translations. While Quest successfully generates multiple translations, its effectiveness as a selection mechanism for a single high-quality response remained unexplored until now. Earlier works have applied MCMC approaches to both autoregressive and masked language models (MLMs). For instance, Miao et al. (2018) and Zhang et al. (2020) use MH with a proposal distribution that makes token-level modifications for constrained generation tasks.
In the case of MLMs, previous works have explored various forms of Gibbs sampling (Berglund et al., 2015; Su et al., 2018; Wang & Cho, 2019; Yamakoshi et al., 2022). However, as Goyal et al. (2022) show, the conditional distributions from MLMs result in invalid Gibbs samplers. In response, they propose an MH correction on the masked conditionals, resulting in higher quality generations. Building on this, Mireshghallah et al. (2022) and Forristal et al. (2023) apply MLMs for controlled text generation. Furthermore, several works have adapted Hamiltonian MCMC algorithms originally designed for high-dimensional continuous distributions (Duane et al., 1987; Neal, 2011) to discrete text generation (Kumar et al., 2022; Qin et al., 2022; Amini et al., 2023; Du et al., 2023).
@@
Test-time scaling methods have emerged as an important approach for improving LM performance without additional training. While our work focuses on local approximations to the optimal aligned distribution from Eq. 2, a parallel line of research explores heuristic search strategies. Best-of-$n$ (BoN; Gao et al., 2022; Stiennon et al., 2020), majority voting (MV; Wang et al., 2023b), weighted majority voting (WMV; Li et al., 2023) are examples of such approaches. As we outline in§ 3.23.3, BoN and WMV can be interpreted as doing test-time alignment.
Furthermore, the RMs we consider in this work are outcome-based RMs, they only work on full generations. Several recent methods use process-based RMs (PRMs) that evaluate partial generations (Lightman et al., 2023; Wang et al., 2024; Uesato et al., 2022). Many techniques have been developed to take advantage of PRMs through guided beam search (Xie et al., 2023), Monte-Carlo tree search (MCTS; Liu et al., 2024; Zhang et al., 2023), and REBASE (Wu et al., 2024). However, as outlined in DeepSeek-AI (2025), PRMs face significant practical limitations.
@@ 6 Conclusion and Future work
Acknowledgments
We thank Alisa Liu and Assaf Harel for their helpful and constructive feedback on the initial versions of the paper. This work was supported in part by NSF grant 211353 and NVIDIA resources provided through the National AI Research Resource Pilot (NAIRR).
@@ A.2 Extreme Value Theory
Following classic results from extreme value theory (Fisher & Tippett, 1928; David & Gumbel, 1960), we know that the distribution of the maximum (appropriately rescaled) converges to a Gumbel distribution. For the specific case where we assume that for a given prompt $x$ the reward function values are distributed as a mixture of two Normal distributions with means $\mu_{1}(x)$ and $\mu_{2}(x)$, variances $\sigma_{1}^{2}(x)$ and $\sigma_{2}^{2}(x)$, and mixture weights $w_{1}(x)$ and $w_{2}(x)$ (where $w_{1}(x)+w_{2}(x)=1$). Appendix C provides empirical evidence that this assumption is reasonable. The tail behavior is dominated by the Gaussian component with the heavier tail, typically the one with larger variance (or larger mean if variances are equal).
@@
The work of Hall (1979) provides asymptotic expressions for location and scale parameters. In particular, one finds that the maximum reward is approximately:
@@ A.3 Distribution of Rewards of $\pi(y\mid x)$
Given the aligned distribution $\pi(y\mid x)$ from Eq. 2, we want an expression for the distribution of rewards under samples from $\pi$. Following the empirical observations from Appendix C, let us assume that for a given prompt $x$, the reward values under the base language model $p_{\text{LM}}(y\mid x)$ follow a mixture of two Normal distributions with means $\mu_{1}(x)$ and $\mu_{2}(x)$, variances $\sigma_{1}^{2}(x)$ and $\sigma_{2}^{2}(x)$, and mixture weights $w_{1}(x)$ and $w_{2}(x)$ (where $w_{1}(x)+w_{2}(x)=1$), i.e.:
@@ A.4 Relating $n$ and $\beta$ via Mode Matching
Appendix B Computational Cost of QAlign

### v2

@@ 3.3 Best-of-$n$ Sampling
This procedure implicitly biases the output distribution toward higher-reward responses. It is easy to see that in the limit of $n$, we are sampling from the maximum reward distribution; however, BoN is remarkably performant even when $n$ is small (fern2022qualityaware; gao2022scalinglawsrewardmodel; llama3).
With some strong assumptions, we can show that the probability density of the BoN distribution can be approximated by our target distribution $\pi(y\mid x)$ from Eq. 1 in the special case when $\beta$ takes the form:
$\beta(n)=\frac{\sigma_{d}(x)}{\sqrt{2\log n_{d}}-\frac{(\log\log n_{d}+\log(4\pi))}{2\sqrt{2\log n_{d}}}}.$ (10)
where $\sigma_{d}(x)$ is the standard deviation of the dominant component of the distribution of reward, under the assumption of Gaussian mixture, which we empirically observe (see Appendix E), when evaluated on samples from our base LM conditioned on $x$, and $n_{d}=w_{d}(x)n$, the dominant component’s sample size, where $w_{d}(x)$ is the weight of the dominant component. A full derivation of this approximation is provided in Appendix A.
This rough equivalence, not noted in past literature to our knowledge, provides a new insight into a key distinction between BoN and the other methods we have discussed, including QAlign. In QAlign, we explicitly define a target distribution $\pi(y\mid x)$, and increasing the test-time budget $n$ refines our approximation of this fixed objective. In contrast, BoN does not converge to a predefined distribution; instead, increasing $n$ progressively shifts the selection process toward responses with higher rewards and deviating further from $p_{\text{LM}}(y\mid x)$. Insofar as the reward model is well-aligned with human preferences, this should make BoN highly effective. However, in realistic settings where the reward is imperfect, increasing $n$ can lead to over-optimization, i.e., selecting responses that score well under the reward model but degrade in actual quality. As a result, BoN can suffer from diminishing returns or even performance degradation as test-time compute increases, which we observe in our experiments (§4.14.2) and previous work (gao2022scalinglawsrewardmodel).
@@ 4.1 Task-Specific Tuning
We evaluate QAlign for task-specific tuning using Llama-3.1-8B-Instruct (llama3) as the base model and a custom bt RM trained on our preference dataset for mathematical reasoning (training details in Appendix H). This RM was initialized with Llama-3.1-8B-Instruct and trained on 64 model-generated response pairs per prompt. We selected pairs based on ground truth answers. [^2] Data and RM will be provided upon publication.
@@
We compare against BoN, and WMV sampling applied to Llama-3.1-8B-Instruct using our trained RM and MV from samples from Llama-3.1-8B-Instruct. For each of the methods we generate $1024$ solutions per problem.
@@
We evaluate on GSM8K (gsm8k), a dataset of grade school math problems, and GSM-Symbolic (gsmsymb), a new benchmark designed to assess out-of-distribution generalization in mathematical reasoning. As demonstrated by gsmsymb, GSM-Symbolic exposes substantial performance degradation across state-of-the-art LLMs exhibiting drops of up to $65\%$ compared to their GSM8K scores when faced with simple variations like changed numerical values and extra redundant clauses.
@@ Results
The results of the task-specific tuning experiments are summarized in Figure 2. QAlign shows progressively lower error rates across both problems as we spend more computational resources. In contrast, on GSM8K, MV shows initial improvement but quickly saturates, while BoN displays improvement with additional compute, temporarily outperforming QAlign, but eventually reaches an inflection point, after which error rates begin to increase. This behavior aligns with the observations from gao2022scalinglawsrewardmodel and theoretical results from §3.3. WMV performs well on GSM8K but becomes relatively worse after a budget of approximately $2^{8}\times 10^{12.51}$ FLOPs.
While all methods show some performance drop on GSM-Symbolic compared to GSM8K, the magnitude varies significantly. BoN and WMV, which rely on the RM, struggle more with the distribution shift, likely because the RM is fit to particular GSM8K-specific patterns. Relative to these two, MV shows greater robustness to these changes. This is primarily because under the distribution shift the RM becomes less reliable, and since MV does not use it, MV’s performance degrades less. However, even under these circumstances, QAlign is able to leverage the RM and outperform MV on this dataset with enough compute. This suggests that QAlign is robust to imperfections in the RM and able to extract a useful signal even when the RM’s reliability is compromised.
@@ 4.2 General Alignment
We evaluate QAlign for more general alignment using Tülu3-8B-SFT as the base model and Tülu3-8B-RM as the RM. The Tülu3 model family (tulu3) was selected for being fully open source, providing instruction-tuned, and aligned versions, along with their corresponding training code and datasets. This enables direct comparison of QAlign on top of the instruction-tuned model to the aligned model, sharing the initial starting point. Additionally, they provide an RM trained on the data used to train the DPO model, which we use in our experiments.
@@
We evaluate again on GSM8K; we add MATH500, a verified subset of the MATH test dataset that contains advanced mathematical problems across algebra, geometry, and calculus (hendrycksmath2021), MMLU-Redux (gema2025mmlu), a refined subset of commonsense and academic knowledge questions, TruthfulQA (truthfulqa), which contains misleading questions designed to elicit truthful responses, and IFEval (ifeval), which measures adherence to complex multi-step instructions. Among these datasets, IFEval is an open-ended generation and, therefore, requires the use of MBR with ROUGE-$1$ instead of MV and WMV. For simplicity, we still refer to the method as MV and WMV to denote MBR from base model samples and the weighted version with importance sampling, respectively, in our comparisons.
@@
For all datasets, we sampled the model with a temperature of 1.0 and prompted it to provide reasoning. Complete prompts are available in Appendix F. As with the mathematical reasoning experiments, we tuned the $\beta$ parameter (in this case, $\beta=0.5$), for Tülu3-8B-RM, using $128$ samples from the GSM8K training data to achieve an average acceptance rate of $50\%$, and using the same $\beta$ for all datasets.
@@ Results
The results of the general alignment experiments are summarized in Table 1. Figure 1 plots the average error rate across all of the datasets as a function of the floating point operations (FLOPS), and Appendix G contains all of the error plots for each individual problem. Similar to our task-specific findings, QAlign, MV, and WMV consistently reduce error rates across all five general alignment datasets as computation increases, with MV exhibiting early saturation. However, similar to the results from GSM-Symbolic, MV outperforms both WMV and BoN. MV applied to the DPO model begins with lower error rates but saturates earlier, ultimately reaching error rates that are worse than QAlign’s final results. Furthermore, we observe that BoN’s inflection point occurs at a lower computational budget than observed in task-specific experiments. This suggests that for real-world alignment problems, BoN is not an effective test-time approach, while QAlign maintains its improvement trajectory even with general RMs.
@@
Table 1: Overview of the results on general alignment. Results show accuracy percentages with fixed 1024 sampled solutions across different methods and models. Our proposed QAlign outperforms other approaches on most benchmarks, achieving the highest scores on all datasets when applied to Tülu3-8B-SFT, and attaining an average performance better than Tülu3-8B-DPO. The highlighted cells indicate best result per benchmark.
@@ 5 Related Work
Our work builds on Quest (quest), which uses MCMC to sample diverse translations. While Quest successfully generates multiple translations, its effectiveness as a selection mechanism for a single high-quality response remained unexplored until now. Earlier works have applied MCMC approaches to both autoregressive and masked language models (MLMs). For instance, cgmh and treesearchcgmh use MH with a proposal distribution that makes token-level modifications for constrained generation tasks.
In the case of MLMs, previous works have explored various forms of Gibbs sampling (berglund2015bidirectional; su2018incorporating; berthasmouth; yamakoshi-etal-2022-probing). However, as goyal2022exposing show, the conditional distributions from MLMs result in invalid Gibbs samplers. In response, they propose an MH correction on the masked conditionals, resulting in higher quality generations. Building on this, mixandmatch and forristal2023block apply MLMs for controlled text generation. Furthermore, several works have adapted Hamiltonian MCMC algorithms originally designed for high-dimensional continuous distributions (hmc; Neal2011ProbabilisticIU) to discrete text generation (kumar2022gradientbased; qin2022cold; amini2023; du2023principled).
@@
Test-time scaling methods have emerged as an important approach for improving LM performance without additional training. While our work focuses on local approximations to the optimal aligned distribution from Eq. 1, a parallel line of research explores heuristic search strategies. Best-of-$n$ (BoN; gao2022scalinglawsrewardmodel; stienon2020learning), majority voting (MV; selfconsistency), weighted majority voting (WMV; li-etal-2023-making) are examples of such approaches. As we outline in§ 3.23.3, BoN and WMV can be interpreted as doing test-time alignment.
The RMs we consider in this work are outcome-based RMs, they only work on full generations. Several recent methods use process-based RMs (PRMs) that evaluate partial generations (lightman2023letsverifystepstep; wang2024mathshepherdverifyreinforcellms; uesato2022solvingmathwordproblems). Many techniques have been developed to take advantage of PRMs through guided beam search (xie2023selfevaluationguidedbeamsearch), Monte-Carlo tree search (MCTS; liu2024dontthrowawayvalue; zhang2023planninglargelanguagemodels), and REBASE (wu2024inferencescalinglawsempirical). However, as outlined in deepseekai2025deepseekr1, PRMs face significant practical limitations.
Furthermore, recent theoretical work has analyzed test-time scaling methods, particularly BoN sampling. farinhas2025rerankinglawslanguagegeneration; schaeffer2025largelanguagemonkeyspower derive BoN scaling laws relating the number of generated samples to performance gains. gui2024bonbonalignmentlargelanguage prove that BoN is optimal for the trade-off between win-rate against the base model and KL divergence from the base model. beirami2025theoreticalguaranteesbestofnalignment provide win-rate guarantees and derive a closed-form probability mass function for the BoN policy, along with a new KL estimator. yang2024asymptoticslanguagemodelalignment show that BoN is asymptotically equivalent to the optimal policy for reward maximization under KL constraints, assuming a memoryless LM and linear reward function. Our work extends this research direction by establishing a novel approximation connecting the BoN distribution with $n$ samples to the optimal aligned distribution with parameter $\beta$ (§ 3.3).
@@ A.2 Extreme Value Theory
Following classic results from extreme value theory (Fisher_Tippett_1928; David1960StatisticsOE), we know that the distribution of the maximum (appropriately rescaled) converges to a Gumbel distribution. For the specific case where we assume that for a given prompt $x$ the reward function values are distributed as a mixture of two Normal distributions with means $\mu_{1}(x)$ and $\mu_{2}(x)$, variances $\sigma_{1}^{2}(x)$ and $\sigma_{2}^{2}(x)$, and mixture weights $w_{1}(x)$ and $w_{2}(x)$ (where $w_{1}(x)+w_{2}(x)=1$). Appendix E provides empirical evidence that this assumption is reasonable. The tail behavior is dominated by the Gaussian component with the heavier tail, typically the one with larger variance (or larger mean if variances are equal).
@@
The work of 1beb852f-e02d-3c6a-88a8-d80908be5528 provides asymptotic expressions for location and scale parameters. In particular, one finds that the maximum reward is approximately:
@@ A.3 Distribution of Rewards of $\pi(y\mid x)$
Given the aligned distribution $\pi(y\mid x)$ from Eq. 1, we want an expression for the distribution of rewards under samples from $\pi$. Following the empirical observations from Appendix E, let us assume that for a given prompt $x$, the reward values under the base language model $p_{\text{LM}}(y\mid x)$ follow a mixture of two Normal distributions with means $\mu_{1}(x)$ and $\mu_{2}(x)$, variances $\sigma_{1}^{2}(x)$ and $\sigma_{2}^{2}(x)$, and mixture weights $w_{1}(x)$ and $w_{2}(x)$ (where $w_{1}(x)+w_{2}(x)=1$), i.e.:
@@ A.4 Relating $n$ and $\beta$ via Mode Matching
Appendix B RLHF as Bayesian Inference
For a fixed prompt $x$, the RLHF objective from Eq. 2, i.e.:
$\mathcal{L}(\theta,x)=\mathbb{E}_{x\sim\mathcal{D}}\big[\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\left[r_{\phi}(y,x)/\beta\right]-D_{\text{KL}}\left(q_{\theta}(y\mid x)\ \|\ p_{\text{LM}}(y\mid x)\right)\big],$
where $q_{\theta}(y\mid x)$ is the learned policy, is maximized by the posterior
$p(y\mid\gamma=1,x)=\frac{1}{Z_{\beta}(x)}\,p_{\text{LM}}(y\mid x)\exp\left({r(y,x)}/{\beta}\right),$
which we denote by $\pi_{\beta}(y\mid x)$.
Proof.
Following the definitions in §  2, $\gamma=1$ denotes a “preferred” response under the human preference model, and $Z_{\beta}(x)$ is the normalizing constant. To prove this, we follow the standard variational inference derivation starting from the log evidence $\log p(\gamma=1\mid x)$.
We begin by writing
$\log p(\gamma=1\mid x)=\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\left[\log p(\gamma=1\mid x)\right].$
By Bayes’ rule we can rewrite the log evidence as
$\log p(\gamma=1\mid x)=\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\left[\log\frac{p(y,\gamma=1\mid x)}{p(y\mid\gamma=1,x)}\right].$
We now multiply and divide inside the logarithm by $q_{\theta}(y\mid x)$:
$\log p(\gamma=1\mid x)=\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\left[\log\frac{p(y,\gamma=1\mid x)\,q_{\theta}(y\mid x)}{p(y\mid\gamma=1,x)\,q_{\theta}(y\mid x)}\right].$
This yields the standard variational decomposition:
$\log p(\gamma=1\mid x)=\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\left[\log\frac{p(y,\gamma=1\mid x)}{q_{\theta}(y\mid x)}\right]+D_{\mathrm{KL}}\!\big(q_{\theta}(y\mid x)\ \|\ p(y\mid\gamma=1,x)\big).$
Next we expand the joint distribution $p(y,\gamma=1\mid x)=p(\gamma=1\mid y,x)\,p_{\text{LM}}(y\mid x)$ which decomposes further into:
$\log p(\gamma=1\mid x)=\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\!\left[\log p(\gamma=1\mid y,x)\right]-D_{\mathrm{KL}}\!\big(q_{\theta}(y\mid x)\ \|\ p_{\text{LM}}(y\mid x)\big)+D_{\mathrm{KL}}\!\big(q_{\theta}(y\mid x)\ \|\ p(y\mid\gamma=1,x)\big).$
We now plug in the preference model (§ 2)
$p(\gamma=1\mid y,x)=\exp\!\left((r(y,x)-\max_{y}r(y,x))/\beta\right),$
and note that both $\max_{y}r(y,x)$ and $\log p(\gamma=1\mid x)$ are constants that depend only on $x$. Thus we can rewrite the previous identity as
$c=\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\!\left[\frac{r(y,x)}{\beta}\right]-D_{\mathrm{KL}}\!\big(q_{\theta}(y\mid x)\ \|\ p_{\text{LM}}(y\mid x)\big)+D_{\mathrm{KL}}\!\big(q_{\theta}(y\mid x)\ \|\ p(y\mid\gamma=1,x)\big),$
for some constant $c$ depending only on $x$.
Rearranging gives
$c-\mathbb{E}_{y\sim q_{\theta}(y\mid x)}\!\left[\frac{r(y,x)}{\beta}\right]+D_{\mathrm{KL}}\!\big(q_{\theta}(y\mid x)\ \|\ p_{\text{LM}}(y\mid x)\big)=D_{\mathrm{KL}}\!\big(q_{\theta}(y\mid x)\ \|\ p(y\mid\gamma=1,x)\big).$
Since the left-hand side is exactly $c-\mathcal{L}(\theta,x)$, minimizing the KL divergence
$D_{\mathrm{KL}}\big(q_{\theta}(y\mid x)\ \|\ p(y\mid\gamma=1,x)\big)$
is equivalent to maximizing the RLHF objective $\mathcal{L}(\theta,x)$. Therefore the maximizer in $q_{\theta}$ is
$q_{\theta}(y\mid x)=p(y\mid\gamma=1,x),$
i.e., the posterior distribution
$p(y\mid\gamma=1,x)=\frac{1}{Z_{\beta}(x)}\,p_{\text{LM}}(y\mid x)\exp\left({r(y,x)}/{\beta}\right).$
This shows that $\pi_{\beta}(y\mid x)$ is the optimal policy for a single prompt.
Appendix C QAlign Convergence proof
Let $\pi_{\beta}(y\mid x)$ be the target distribution, and consider the Metropolis–Hastings chain used by QAlign with target $\pi_{\beta}$ and the Quest proposal kernel $q(y\mid y^{t},x)$. Then the distribution of $y^{t}$ converges in distribution to $\pi_{\beta}(y\mid x)$ as $t\to\infty$.
Proof.
The theory comes from the standard Metropolis–Hastings (MH; mh) together with the properties of the proposal distribution introduced in Quest (quest). Once we specify (i) the target distribution $\pi_{\beta}(y\mid x)$ and (ii) a valid proposal kernel, the MH theorem ensures that the resulting Markov chain converges to $\pi_{\beta}$ in the limit.
The central requirement is that the MH transition kernel satisfies detailed balance with respect to $\pi_{\beta}$:
$\pi_{\beta}(y^{t}\mid x)\,T(y^{t}\rightarrow y)=\pi_{\beta}(y\mid x)\,T(y\rightarrow y^{t}),$
where $T(\cdot\to\cdot)$ denotes the one-step transition probability of the chain. For the standard MH acceptance rule,
$\alpha_{\beta}(y,y^{t})=\min\!\left(1,\frac{\pi_{\beta}(y\mid x)\,q(y^{t}\mid y,x)}{\pi_{\beta}(y^{t}\mid x)\,q(y\mid y^{t},x)}\right),$
mh shows that detailed balance holds automatically for any proposal $q(y\mid y^{t},x)$ that is irreducible and aperiodic.
The work of Quest (quest) proves that the proposal distribution we use, defined using LLM-based random suffix resampling, is irreducible and aperiodic. An irreducible Markov chain is one that for any two responses $y,y^{\prime}$ in the support of the LM, the proposal kernel $q(y^{\prime}\mid y,x)$ has non-zero probability of reaching $y^{\prime}$ from $y$ in a finite number of proposal steps. In our case, this is ensured by the fact that the proposal mechanism can always select index 0 and regenerate an entire response from scratch.
An aperiodic Markov chain is one in which every state is aperiodic (equivalently, every state has period 1). A state is periodic if returns to that state can occur only after numbers of steps that are all multiples of some integer greater than 1, for example only after an even number of steps, or only after multiples of 3. Otherwise, the state is aperiodic, meaning there is no such mandatory multiple. A simple sufficient condition for aperiodicity is that each state has a non-zero probability of returning to itself in one transition (a self-loop). In our setting, this holds because with positive probability we sample index 0 (which occurs with probability $1/|y|$) and regenerate the entire response, and this regeneration can reproduce the current response with non-zero probability, giving a one-step self-return.
Thus, combining detailed balance with irreducibility and aperiodicity, the Markov chain ergodic theorem(Neal2011ProbabilisticIU) guarantees that the chain convergence to the optimal aligned distribution $\pi_{\beta}$ in the limit $t\rightarrow\infty$.
Appendix D Computational Cost of QAlign